find the area of an isosceles right triangle whose one of the equal sides is 10 cm long.
Answer by ewatrrr(24785)

isosceles right triangle: the sides are in the ratio 1:1: √2
When one of the equal sides is 10 ...b = 10√2 and (1/2)base = 5√2
h = sqrt(100 - 50) Pythagorean Theorem
h = 5√2
A = (1/2)bh = (1/2)(10√2)(5√2) = 50 cm^2